OpenStudy (anonymous):

Perform the necessary operations. -5-9+(-4)-8=

OpenStudy (anonymous):

the positive and negative make it a negative and then we just subtract the rest:\[-5-9+(-4)-8=\]\[-5-9-4-8=\]\[-14-4-8=\]\[-18-8=-26\]

OpenStudy (anonymous):

If you know the order of operations, this shouldn't be too tricky of a problem. In the order of operations, addition and subtraction are in the same step. This means that the best approach would be to solve the problem straight through from left to right because it doesn't matter the order of addition and subtraction.

OpenStudy (anonymous):

I see... I copied and pasted these responses to my notes!! Thanks yall!!

OpenStudy (anonymous):

welcome :)

OpenStudy (anonymous):

Good idea! You're welcome.